According to a CBS/New York Times poll taken in 1992, 15% of the public have responded to a telephone
call-in poll. In a random group of five people, what is the probability that exactly two have responded to
a call-in poll? - correct answer binompdf(5, .15, 2) = 0.1382
According to a medical study, 40% of people will experience an adverse side effect within one hour after
taking a particular experimental drug. Suppose you survey 15 people who participated in the study.
What is the probability that at least 8 of them experienced an adverse effect? - correct answer
P(X≥8)= 1 - binomcdf(15, 0.4, 7) = 0.2131
In a certain mechanical system, 5 bulbs are used and will operate if at least three are functioning
properly. Given that 15% of bulbs are defective, what is the probability that the system will malfunction?
- correct answer P(X<3) = binomcdf(5, .15, 2)= .9734
An automobile-safety researcher, Kathryn, claims that 1 in 10 automobile accidents are caused by driver
fatigue. What is the probability that at most three out of five automobile accidents are caused by driver
fatigue? - correct answer P(X≤3)=binomcdf(5, .10, 3)=.9995
If you had no idea how to do any of the problems on a 10 question multiple choice quiz and you chose to
randomly select your answer, what is the probability you would get 70% of them correct?
(assume there are 4 choices per question). - correct answer P(X=7) =
binompdf(10, .24, 7) = 0.003
If you had no idea how to do any of the problems on a 10 question multiple choice quiz and you chose to
randomly select your answer, what would the probability be that you got a 70% or better on the exam? -
correct answer P(X ≥7)= 1 - binomcdf(10, .25, 6) = .004
A pet food manufacturer is promoting a new brand with a rebate offer on its 10-pound bag. Each
package is supposed to contain a coupon for a $4.00 mail-in rebate. The company has found that the
machine dispensing these coupons fails to place a coupon in 10 percent of the bags. If a dog owner buys
5 bags, find the probability that at least 1 bag will fail to have a coupon. - correct answer
P(X≥1)= 1 - binompdf(5, .10, 0) = 0.4095